OpenStudy (anonymous):

simplify the square root of 8x^3

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{8x^3}\] \[\Large \sqrt{4*2*x^2*x}\] \[\Large \sqrt{4}*\sqrt{2}*\sqrt{x^2}*\sqrt{x}\] \[\Large 2*\sqrt{2}*x*\sqrt{x}\] \[\Large 2x\sqrt{2x}\] So \[\Large \sqrt{8x^3}=2x\sqrt{2x}\] for all nonnegative values of x.

OpenStudy (ash2326):

We have \[\sqrt{8x^3}\] We know \[8=2\times 2 \times 2\] and \[x^3=x \times x \times x\] so we have \[\sqrt {8x^3}=\sqrt{2\times 2 \times 2 \times x \times x \times x}\] or \[\sqrt{2^2 \times 2 \times x^2 \times x}\] so we get \[2x\sqrt{2x}\]